OpenStudy (anonymous):

This morning, Robyn wrote a check in the amount of $35.80. This brought her balance down to –$18.25. Suppose n represents Robyn's balance before she wrote the check. Which equation can be used to represent this situation? A. n – 18.25 = 35.80 B. n – 35.80 = –18.25 C. n + 35.80 = –18.25 D. n – 18.25 = –35.80

OpenStudy (anonymous):

This morning, Elena wrote a check in the amount of $42.50. This brought her balance down to –$14.25. Suppose n represents Elena's balance before she wrote the check. Which equation can be used to represent this situation? A. n – 14.25 = 42.50 B. n + 42.50 = –14.25 C. n – 42.50 = –14.25 D. n – 14.25 = –42.50

OpenStudy (anonymous):

i coshe b on #1. i chose b one #2

OpenStudy (anonymous):

#1 b #2 c sorry

OpenStudy (pandagurl):

Both are correct :D Good Job!